OpenStudy (anonymous):

There are 5 red and 9 green balls in a bag. If you randomly choose balls one at a time, with replacement, what is the probability of choosing 2 red balls and then 1 green ball? My answer was 225/2744. am i right????!!!

Parth (parthkohli):

You are replacing....so it'd be P(red) * P(red) * P(green)

OpenStudy (anonymous):

Here are the choices.. sorry forgot to include them :/

OpenStudy (anonymous):

I got (5/14)(5/14)(9/14)=225/2744... is that right? or no?? I am confused..

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow {5 \times 5 \times 9 \over 14^3 } }\) \(\Large \color{Black}{\Rightarrow {225 \over 2744} }\) Correct!